The profit for a product is given by ​P(x)equals=negative 11 x squared plus 1320 x minus 38 comma 500−11x2+1320x−38,500​, where x is the number of units produced and sold. How many units give break even​ (that is, give zero​ profit) for this​ product?
Answer #1

4FD068A2-D5EB-4453-B3B6-B655348E6671.png

source: Myself
answered by: Kp Rajput
Answer #2

P(z) =-11x2 + 1 320x-38500

for break even take P(x)=0

−11x^2+1320x−38500=0−11x^2+1320x−38500=011 1320r - 38500 0

-11 (22-1202. +3500) = 0

2-120x 35000

x^2-50x-70x+3500=0

r(z-50)-70(-50) = 0

(z-50) (2-70) = 0

50. 2. 70

when 50 and 70 units are sold then break even occurs